# Fine-Tuning Image-Conditional Diffusion Models is Easier than You Think

### v1

@@ Appendix A Implementation and Evaluation Details
Since Virtual KITTI 2 does not provide annotations for surface normals, we compute them ourselves with the ground-truth depth maps, employing discontinuity-aware gradient filters from [15]. A qualitative example of the resulting normals can be seen in Fig. 7.
@@
Figure 7: Virtual KITTI 2 example. Top: Synthetic RGB image. Middle: Ground-truth depth map. Bottom: Ground-truth surface normals, generated using discontinuity-aware gradient filters [15].
@@
For most existing methods in Sec. 6 and Sec. 6 we obtain the performance metrics either from the papers introducing these methods or from the Marigold and DSINE papers. The missing scores, like those of the newer GeoWizard [16] and DepthFM [18] models, are obtained by reevaluating the respective models with their official inference code and released checkpoints. In the case of DepthFM, the prediction alignment with respect to the ground-truth metric depth happens in the log metric space.
Figure 8: Single-step outputs of Marigold and Stable Diffusion. With a single step, Stable Diffusion produces a blurry image at best, while Marigold outputs a sensible depth map. Note that the input prompt is text for Stable Diffusion, but an RGB image for Marigold.
@@
Figure 9: Few-step inference of Marigold and Stable Diffusion. With more steps, the adverse effects of the broken DDIM scheduler get less pronounced. Both Marigold and Stable Diffusion produce sharper outputs with more steps, but the difference is much greater for Stable Diffusion.
Appendix B DDIM Inference
During training, the highest noise level corresponds to the last timestep $t=T$, and $t=1$ corresponds to a very small noise level. The DDIM inference scheduler iterates over a series of $k$ timesteps $\tau_{1}>\tau_{2}>\ldots>\tau_{k}>0$ and iteratively denoises the initial noise input $\bm{\mathbf{z}}_{\tau_{1}}$. We consider the leading and trailing schedules that are also discussed by Lin et al. [30] and show the selected timesteps for different $k$ in LABEL:tab:leading_vs_trailing. The original leading timestep selection strategy of the DDIM scheduler excludes the final timestep $T$. This leads to a mismatch between training and inference; using the leading schedule, the model expects a partially denoised input in the first step, but instead receives only noise. In contrast, the fixed trailing strategy always starts with $t=T$ for the first denoising step, properly aligning training and inference. In the limit of $k\rightarrow T$ inference steps, both strategies converge to the same behavior.
In Fig. 8, we illustrate the difference between single-step predictions using the broken leading and the fixed trailing DDIM scheduler for Marigold [26] and Stable Diffusion [41]. Both models output noise when using the broken scheduler. With the fixed implementation, both models predict the mean of their respective conditional distribution. For single-step Marigold this results in a well-defined depth map, whereas for single-step Stable Diffusion, it produces a blurry image with coarse structures that roughly align with the input prompt.
Fig. 9 further demonstrates the scheduler’s impact when multiple steps are considered. It clearly shows that the effect of the broken scheduler becomes less noticeable as the number of inference steps increases. Additionally, the weak text conditioning in Stable Diffusion leads to blurry images, which gradually sharpen as more inference steps are taken. In contrast, the strong image conditioning in Marigold allows the model to predict reasonably accurate depth maps already in the first step. As shown by the heatmap in Fig. 2, subsequent steps only lead to small changes in the predicted distances, and most of the scene remains unchanged.
Table 10: Frozen vs. fine-tuned VAE decoder. We conduct end-to-end fine-tuning of Marigold [26] for depth estimation, and assess the effect of freezing or fine-tuning the weights of the pretrained VAE decoder.
Decoder NYUv2 [46] KITTI [17] ETH3D [45] ScanNet [9] DIODE [48]
Frozen 5.2 96.6 9.6 91.9 6.2 95.9 5.8 96.2 30.2 77.9
Fine-tuned 5.3 96.5 9.6 91.9 6.2 96.0 5.8 96.1 30.2 77.7
Appendix C Additional Experimental Results
@@
[…]
@@
DepthFM [18] proposes a direct mapping from input images to depth maps through flow matching, leveraging Stable Diffusion v2 [41] as a prior. We observe that, apart from the ETH3D $\delta$1 and DIODE [48] metrics, a simpler approach like E2E FT achieves better performance with a more than $10\times$ speedup as seen in Tab. 9.
@@
By default, we keep the pretrained VAE decoder frozen while conducting end-to-end fine-tuning. Tab. 10 shows that fine-tuning the weights of this decoder does not improve performance.
Appendix D More Qualitative Samples
In Fig. 10 and Fig. 11, we show additional qualitative results for depth and normal estimation, respectively, comparing Marigold [26] and the end-to-end fine-tuned models. Noticeable over-sharpening and high-frequency noise artifacts appear in the multi-step Marigold predictions, particularly in the normal estimations. Although ensembling techniques help reduce these artifacts, grainy textures can still persist. In contrast, the end-to-end fine-tuned models do not exhibit these issues.
@@ Appendix D More Qualitative Samples
Figure 10: Additional qualitative samples for depth estimation. “Marigold ($X$, YY)” denotes Marigold using $X$ inference steps with an ensemble of size $Y$.
@@
Figure 11: Additional qualitative samples for normal estimation. “Marigold ($X$, YY)” denotes Marigold using $X$ inference steps with an ensemble of size $Y$.

### v2

@@ Appendix A DDIM Inference
In Fig. A-1, we illustrate the difference between single-step predictions using the broken leading and the fixed trailing DDIM scheduler for Marigold [24] and Stable Diffusion [39]. Both models output noise when using the broken scheduler. With the fixed implementation, both models predict the mean of their respective conditional distribution. For single-step Marigold this results in a well-defined depth map, whereas for single-step Stable Diffusion, it produces a blurry image with coarse structures that roughly align with the input prompt.
Fig. A-3 further demonstrates the scheduler’s impact when multiple steps are considered. It clearly shows that the effect of the broken scheduler becomes less noticeable as the number of inference steps increases. Additionally, the weak text conditioning in Stable Diffusion leads to blurry images, which gradually sharpen as more inference steps are taken. In contrast, the strong image conditioning in Marigold allows the model to predict reasonably accurate depth maps already in the first step. As shown by the heatmap in Fig. 2b in the main text, subsequent steps only lead to small changes in the predicted distances, and most of the scene remains unchanged.
Figure A-1: Single-step outputs of Marigold and Stable Diffusion. With a single step, Stable Diffusion produces a blurry image at best, while Marigold outputs a sensible depth map. Note that the input prompt is text for Stable Diffusion, but an RGB image for Marigold.
Appendix B Detailed Experimental Setup
@@
Figure A-2: Virtual KITTI 2 example. Top: Synthetic RGB image. Middle: Ground-truth depth map. Bottom: Ground-truth surface normals, generated using discontinuity-aware gradient filters [13].
@@
Since Virtual KITTI 2 does not provide annotations for surface normals, we compute them ourselves with the ground-truth depth maps, employing discontinuity-aware gradient filters from [13]. A qualitative example of the resulting normals can be seen in Fig. A-2.
@@
For most existing methods in Tab. 5 and Tab. 6 we obtain the performance metrics either from the papers introducing these methods or from the Marigold and DSINE papers. The missing scores, like those of the newer GeoWizard [14] and DepthFM [16] models, are obtained by reevaluating the respective models with their official inference code and released checkpoints. In the case of DepthFM, the prediction alignment with respect to the ground-truth metric depth happens in the log metric space.
Table A-2: Frozen vs. fine-tuned VAE decoder. We conduct end-to-end fine-tuning of Marigold [24] for depth estimation, and assess the effect of freezing or fine-tuning the weights of the pretrained VAE decoder.
Decoder NYUv2 [44] KITTI [15] ETH3D [43] ScanNet [7] DIODE [46]
Frozen 5.2 96.6 9.6 91.9 6.2 95.9 5.8 96.2 30.2 77.9
Fine-tuned 5.3 96.5 9.6 91.9 6.2 96.0 5.8 96.1 30.2 77.7
@@
Figure A-3: Few-step inference of Marigold and Stable Diffusion. With more steps, the adverse effects of the broken DDIM scheduler get less pronounced. Both Marigold and Stable Diffusion produce sharper outputs with more steps, but the difference is much greater for Stable Diffusion.
Appendix C Additional Results
@@
GeoWizard [14] jointly predicts depth and surface normals, using a similar training and evaluation setup as Marigold. We find that GeoWizard suffers from the same flaw in the DDIM implementation as Marigold, and end-to-end fine-tuning the model for depth and normal estimation significantly boosts the performance (see Tab. A-3 and Tab. 3 in the main text). In particular, the fine-tuned model performs better than both the fixed single-step model and the previously best reported results with 50 steps and ensembling of 10 predictions.
@@
DepthFM [16] proposes a direct mapping from input images to depth maps through flow matching, leveraging Stable Diffusion v2 [39] as a prior. We observe that, apart from the ETH3D $\delta$1 and DIODE [46] metrics, a simpler approach like E2E FT achieves better performance with a more than $10\times$ speedup as seen in Tab. A-4.
@@
By default, we keep the pretrained VAE decoder frozen while conducting end-to-end fine-tuning. Tab. A-2 shows that fine-tuning the weights of this decoder does not improve performance.
Further Qualitative Samples.
Fig. A-4 and Fig. A-5 show qualitative results for depth and normals estimation, respectively, comparing Marigold [24] and the end-to-end fine-tuned models. The fixed single-step model fails to produce sharp results, while the multi-step model exhibits noticeable over-sharpening and high-frequency noise artifacts (even after ensembling), particularly in the normals estimations. In contrast, the end-to-end fine-tuned models do not exhibit these issues.
Addendum
We were made aware of recent work by Xu et al. [49]. Similar to us, they directly fine-tune Stable Diffusion in an end-to-end fashion, however, we arrive to this point in a very different way. We initially discovered the issue with the DDIM scheduler, fixed this in Marigold, and in turn arrived to an end-to-end fine-tuning scheme that works for Marigold. Surprisingly, our ablations showed that this also works well for direct fine-tuning of Stable Diffusion. The main contribution of Xu et al. is an approach to fine-tune Stable Diffusion (for a broader spectrum of tasks). However, even with additional modules on top, their method achieves lower scores than some of the baselines. As such, these results might lead one to conclude that end-to-end fine-tuning is not a suitable alternative to multi-step, diffusion-based depth and normal estimation. In contrast, our simple end-to-end fine-tuning setup does outperform diffusion baselines, demonstrating that it is an effective and efficient alternative.
@@
Figure A-4: Additional qualitative samples for depth estimation. “Marigold ($X$, YY)” denotes Marigold using $X$ inference steps with an ensemble of size $Y$.
@@
Figure A-5: Additional qualitative samples for normal estimation. “Marigold ($X$, YY)” denotes Marigold using $X$ inference steps with an ensemble of size $Y$.